During the four years that Mrs Lopez owned her car, she found that her total car expenses were $18000. Fuel and maintenance costs accounted for 1/3 of the total and depreciation accounted for 3/5 of the remainder. The cost of insurance was 3 times the cost of financing, and together these two costs accounted for 1/5 of the total. If the only expenses were taxes and licence fees, then the total cost of financing was how much more or less than the cost of taxes and licence fees?

$300 less is supposed to be the right answer, but I am getting $600 less. Can you please explain...
Thank you

Answer by solver91311(24713)



Fuel/Maint is one-third of total: 1/3 times 18000 = 6000
Depreciation is 3/5 of remainder: 3/5 times 12000 = 7200
Finance/Insurance 1/5 of total: 1/5 times 18000 = 3600

6000 + 7200 + 3600 = 16800

18000 - 16800 = 1200

Ins = 3 times Fin, so F + I = 4 F = 3600, so F = 900 which is 300 less than 1200



John

My calculator said it, I believe it, that settles it